Puzzle of the day · No. 9 · 21 October
A stubborn bike
Lucky’s old motorbike starts on the first try 80% of the time, and every further try has the same chance. What is the chance that Lucky needs more than two tries?
Answer: 4%. More than two tries means the first two failed. Each fails with a chance of 0.2, both in a row 0.2 × 0.2 = 0.04.
Discussion